# Kinodynamic Task and Motion Planning using VLM-guided and Interleaved Sampling

### v1

@@ V-C Real-World Demonstration
We further validated our approach in the Blocksworld domain through a real-world robotic demonstration. The setup consisted of dual UR5e manipulators equipped with Robotiq 3F grippers and an Intel RealSense D455 sensor. Given an RGB image of the tabletop scene captured by the sensor, an open-vocabulary object detection model [49] generated segmented masks for each object. These 2D masks were projected into 3D points using depth values, from which each object was localized in the simulator. We utilized the average CIELAB color value of each object’s point cloud to distinguish individual blocks. Since the geometric and inertial properties of the objects were assumed to be known a priori, planning was performed as described in Sec. IV. The resulting plan, consisting of the arm trajectories and gripper open/close actions, was directly executed on the real robot, as shown in Fig. 4.
We conducted five trials for each $3\leq n\leq 6$ in the Blocksworld domain. The planning times and success rates are similar to the simulation results presented in the previous section. In particular,
@@
VI CONCLUSIONS AND FUTURE WORK
This paper presents a novel kinodynamic TAMP algorithm that interleaves symbolic task planning and continuous action parameter refinement at every hybrid state tree expansion step. Furthermore, a VLM guides the replanning process by selecting promising actions in the hybrid state tree. Although our approach demonstrates strong performance, there are several limitations. The overhead due to physics simulation increases with plan length, and planning performance can be sensitive to the quality of the underlying samplers. In addition, the assumptions of full observability and a black-box transition model are restrictive. Future work will focus on extending our approach beyond pick-and-place style manipulation to more diverse domains, relaxing assumptions such as full observability, and conducting more extensive experiments on physical robots dealing with possible sensing discrepancy issues.

### v2

@@ V-C Real-World Demonstration
We further validated our approach in Blocksworld through a real-world robotic demonstration. The setup consisted of dual UR5e manipulators equipped with Robotiq 3F grippers and an Intel RealSense D455 sensor. Given an RGB image of the tabletop scene captured by the sensor, an open-vocabulary object detection model [45] generated segmented masks for each object. These 2D masks were projected into 3D points using depth values, from which each object was localized in the simulator. Since the geometric and inertial properties of the objects were assumed to be known a priori, planning was performed as described in Sec. IV. The resulting plan, consisting of the arm trajectories and gripper open/close actions, was directly executed on the real robot, as shown in Fig. 4.
We conducted 5 trials for each $3\leq n\leq 6$ in the Blocksworld domain. The planning times and success rates are similar to the simulation results presented in Sec. V. In particular,
@@
VI CONCLUSION AND FUTURE WORK
This paper presents a novel kinodynamic TAMP planner that interleaves symbolic planning with kinodynamic feasibility checks using a physics simulator as the world transition model, while a VLM guides forward search and replanning.
Limitations: A major limitation of this work is that the computational overhead introduced by physics simulation increases with plan length. Another limitation is that performance can be sensitive to the quality of the underlying samplers and variations in the VLM, such as different models, temperature settings, prompt design, and the number of input images. In addition, the assumptions of full observability and a black-box transition model remain restrictive.
Future Work: Future work will focus on extending our approach beyond pick-and-place-style manipulation to more diverse domains such as tool use, deformable objects, and contact-rich tasks. These domains require more expressive abstractions and validation strategies, and could be evaluated on more realistic benchmarks such as [46]. We can also reduce sensitivity to sampler quality by integrating learned sampling strategies to improve efficiency, as well as to relax assumptions such as full observability.
ACKNOWLEDGMENT
This work was supported in part by the ITRC/IITP Program (IITP-2026-RS-2020-II201460), and in part by the NRF (NRF-2022R1A2B5B03001385) in South Korea.